OpenStudy (anonymous):

How do I go about solving the integration of Sqrt(x) (x^2+10) all over x^4 dx?

OpenStudy (anonymous):

deezil you are missing square root of x

OpenStudy (anonymous):

oh sorry i sure am lol lemme fix it

OpenStudy (anonymous):

\[(x^{1/2})(x^2+10)* x ^{-4}\]

OpenStudy (anonymous):

u want answer??

OpenStudy (anonymous):

Okay I've done that and integrated but it doesn't appear right

OpenStudy (anonymous):

what'd you get?

OpenStudy (anonymous):

-2^\[-2^(3/2) (x^3/3 +10x)/9x^3 + c\]

OpenStudy (anonymous):

that 3/2 is an exponent

OpenStudy (anonymous):

yeah i got you, use { } for exponents instead of ( )

OpenStudy (anonymous):

kk thanks, how did you get that?

OpenStudy (anonymous):

what are you talking about division and multiplication rule... he's integrating? Either i'm really out of it or I have no idea what you are talking about

OpenStudy (anonymous):

oh sh&t i've been deriving the whole time! so sorry lemme change it

OpenStudy (anonymous):

I believe you're thinking derivative, this is integration haha

OpenStudy (anonymous):

There is no quotient rule for integration. It is called Integration by parts and that would not be hard at all as you could use tabulars method

OpenStudy (anonymous):

the top goes to 0

OpenStudy (anonymous):

thanks outkast, I have a bad habit of not reading the entire question!

OpenStudy (anonymous):

it's ok i was like what in the world...

OpenStudy (anonymous):

Yeah now i see why our answers were so far off! Well, if you ever need to derive same thing... lol

OpenStudy (anonymous):

ok now I'm getting...

OpenStudy (anonymous):

i believe you would tackle it the same though i can easily do by parts but i'm not sure if he's in calculus 2 or 3

OpenStudy (anonymous):

or 1

OpenStudy (anonymous):

2

OpenStudy (anonymous):

\[-(2(x^2+2))/x^{5/2}\]

OpenStudy (anonymous):

why does it become \[x^2+2\]

OpenStudy (anonymous):

scratch that you cannot use parts since u does not go to 0 i forgot there was a sqrt

OpenStudy (anonymous):

split x^2+10 into 2 integrals

OpenStudy (anonymous):

or simply do it there is something they do at the end but is your answer this \[\frac{-2}{x^{1/2}}-\frac{4}{x^{5/2}}\]

OpenStudy (anonymous):

ahh they use the LCD which is x^2

OpenStudy (anonymous):

I got \[-2x^{3/2}(x^2+10)/9x^3\]

OpenStudy (anonymous):

alright so lets go through this by steps first you want to get everything ready to integrate so times the dtop by x^-4 correct? \[\sqrt{x}(x^2+10)(x^{-4})\] =\[(x^{3/2}+10x^{1/2})(x^{-4}) = x^{-5/2}+10x^{-7/2}\]

OpenStudy (anonymous):

no divide

OpenStudy (anonymous):

well yes multiply by x^-4 my bad

OpenStudy (anonymous):

it was divide by x^4 so you are correct

OpenStudy (anonymous):

now using the integration rule which is add one to the exponent and divide by that number\[\int\limits_{}^{}x^{-3/2}+10x^{-7/2}=-2x^{-1/2}-4x^{-5/2}\]

OpenStudy (anonymous):

yes that is correct i made a mistake in the above it should be \[x^{-3/2} instead of x^{-5/2}\]

OpenStudy (anonymous):

how is it to the -3/2?

OpenStudy (anonymous):

\[x^{5/2}*x^{-8/2} = x^{-3/2}\]

OpenStudy (anonymous):

\[x^{-8/2}=x^{-4}\]

OpenStudy (anonymous):

isn't it \[x^{3/2} * x^{-8/2}\]

OpenStudy (anonymous):

anyways what you have above is \[\frac{-2}{x^{1/2}}-\frac{4}{x^{5/2}}\]

OpenStudy (anonymous):

no because\[x^{1/2}*x^2=x^{1/2}*x^{4/2}\]

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

when dealing with exponents you add correct? so you wouldn't add \[\frac{1}{2}+2=\frac{3}{2}\]

OpenStudy (anonymous):

that is incorrect you'll have to find a lcd and then add

OpenStudy (anonymous):

when dealing with the multiplication of two exponents

OpenStudy (anonymous):

anyways you'll get the asnwer above and then you'll have to get the LCD for the two below which is x^2 \[\frac{(-2)(x^2)-4}{x^{5/2}}=\frac{-2x^2-4}{x^{5/2}}=\frac{-2(x^2+2)}{x^{5/2}}\]

OpenStudy (anonymous):

which is your answer

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem